OpenStudy (anonymous):

Pastaroni is a local Italian restaurant that serves a specialized pasta salad. Included in this salad is macaroni, linguine and bow-tie pasta. The salad contains six times as much linguine as it does macaroni. Both macaroni and linguine sell for $4; bow-tie pasta sells for $8. If the restaurant makes 140 pounds of pasta every day, how many pounds of each type of pasta are needed if the salad sells for $6? This is the last question I need help with. I just need to find the system of equations for this one too. I know how to solve it so yeah, all I need is system of equation.

OpenStudy (anonymous):

@knock Could you help me with this problem also? This is the last one I need help with so yeah.

OpenStudy (anonymous):

Ok we call z=linguine, x=macaroni and y=bow-tie, six times more linguini than macaroni so linguini=6x maybe this is the most difficult part

OpenStudy (anonymous):

Now we have 140 pounds so 6x+x+y=140

OpenStudy (anonymous):

Yeah I was stuck on the 6x linguini thing.

OpenStudy (anonymous):

And 6x*4+x*4+y*8=6 like on the other problem

OpenStudy (anonymous):

I think that's the system

OpenStudy (anonymous):

So 6x+y+z=140 24x+4x+8y=6?

OpenStudy (anonymous):

No, 6x+y+x=140 24x+4x+8y=6*140 (I have bad conection and din't put de *140 above notice that)

OpenStudy (anonymous):

oh okay. so it's 840 at the end. Okay :) Tyvm @knock

OpenStudy (anonymous):

Oh and I noticed it's your 1st day on OpenStudy, so welcome to OS! :)

OpenStudy (anonymous):

yw ^^; do you understand why 6*140 right?

OpenStudy (anonymous):

Yes, the 140 pounds of pasta * $6 (how much salad sells)

OpenStudy (anonymous):

Indeed ^^

OpenStudy (anonymous):

I know how to do that, it's just those 2 word problems confused me a little unlike the others I did. Lol I actually know the stuff unlike some people who just come here for answers.

OpenStudy (anonymous):

Yes, well I hope we can all learn everyday a little and teach everyday a little to each others, altought some people just want the answers, a lot more wants to learn like you and me ^^ So little by little we can build this community between all

OpenStudy (anonymous):

Yep :).

OpenStudy (anonymous):

Agent, there has to be 3 variables I believe, but in his equation there was only 2. Is this correct?

OpenStudy (agent47):

S=6m+m+bow-tie-pasta

OpenStudy (agent47):

Wait lol wrong

OpenStudy (agent47):

Pastaroni is a local Italian restaurant that serves a specialized pasta salad. Included in this salad is macaroni, linguine and bow-tie pasta. The salad contains six times as much linguine as it does macaroni. Both macaroni and linguine sell for $4; bow-tie pasta sells for $8. If the restaurant makes 140 pounds of pasta every day, how many pounds of each type of pasta are needed if the salad sells for $6? S=m+l+b l=6m m+l+b=140

OpenStudy (agent47):

Both macaroni and linguine sell for $4 and pasta is for $8, and the whole salad must be $6 so: 4m+4l+8b=6*140

OpenStudy (anonymous):

K. 4m+4l+8b=840 I'm guessing the 2nd equation is 6m+l+b=140

OpenStudy (agent47):

Wait lol I'm not done yet..l=6m m+l+b=140, m+6m+b=140, or 7m+b=140 4m+4l+8b=840, 4m+24m+8b=840, or 28m+8b=840

OpenStudy (agent47):

So you have: 7m+b=140 7m+2b=210 subtract the 2nd equation from first: -b=-70

OpenStudy (anonymous):

K. That would make b=70

OpenStudy (agent47):

7m+70=140 7m=70 m=10

OpenStudy (agent47):

m+l+b=140 10+l+70=140 l=140-80=60

OpenStudy (anonymous):

Tyvm @ Agent47. I was stuck on this one quite a while and didn't want to get it wrong :)

OpenStudy (anonymous):

Yes Loujoelou the other eqaution is z=linguine=6x, but with the system of two equations and two variables, once you know x you will know z

OpenStudy (anonymous):

It was implicit in the system so it would be easier, the solutions must be z=linguine=60, x=macaroni=10 and y=bow-tie=70

OpenStudy (anonymous):

So it isn't wrong @Agent47

OpenStudy (anonymous):

Do you understand @Loujoelou?

OpenStudy (anonymous):

Yep